OpenStudy (vishweshshrimali5):

\[sin^2 \theta = 1\] Find general solution of \(\theta\). I solved for sin \(\theta\) = \(\pm\)1 and got \(\theta = n\pi + (-1)^n \cfrac{\pi}{2}\) or \(\theta = n\pi + (-1)^n \cfrac{3\pi}{2}\) How to find combined solution

OpenStudy (lgbasallote):

sometimes....i doubt you're really in high school.....

OpenStudy (vishweshshrimali5):

Why ?

OpenStudy (lgbasallote):

that kind of trigonometry is rarely taught in high school...

OpenStudy (lgbasallote):

...or maybe i just don't know what that is...

OpenStudy (vishweshshrimali5):

That is general solution to trigonometric equation

OpenStudy (vishweshshrimali5):

@estudier @experimentX

OpenStudy (shubhamsrg):

why dont you treat it like this: 1 = sin^2 @ or cos^2 @ = 0 easier ?

OpenStudy (vishweshshrimali5):

then I have cos\(\theta\) = 2n\(\pi\) \(\pm\) \(\cfrac{\pi}{2}\)

OpenStudy (shubhamsrg):

you mean @ = that,, yep//

OpenStudy (vishweshshrimali5):

Yes